OpenStudy (anonymous):

What is the first term of an arithmetic sequence with a common difference of -7 and a seventh term of 40

OpenStudy (anonymous):

\[a_n = a+d(n-1) \] an= term value = 40 d = -7 n = 7 a = first term

OpenStudy (anonymous):

an=a+(n-1)d 40=a+(7-1)(-7) 40=a-42 82=a. an is the term you want to know a is the first term n is the term number d is the difference

OpenStudy (anonymous):

i mean an is "any term" so in this case its the seventh term number